At what temperature will salt solution boil if 20 g salt (NaCl) is added to 1.5 kg of water?

Note: Remember that NaCl is a strong electrolyte and dissociates to Na+ and Cl- , respectively.


1
Expert's answer
2022-04-14T07:46:04-0400

Solution:

A solution will boil at a higher temperature than the pure solvent.

This is the colligative property called boiling point elevation.

An equation has been developed for this behaviour:

Δt = i × Kb × m

where:

Δt = the change in boiling point

i = the van't Hoff factor

Kb = the molal boiling-point elevation constant

m = the molality of the solute


The van't Hoff factor is the number of particles (ions) formed by the compound in the solution.

The dissociation of NaCl is given by:

NaCl → Na+ + Cl¯

Therefore, van’t Hoff factor for NaCl is 2


Kb for water is 0.512°C kg mol¯1


The molar mass of NaCl is 58.44 g mol¯1

Therefore,

Moles of NaCl = (20 g NaCl) × (1 mol NaCl / 58.44 g NaCl) = 0.342 mol


Molality of NaCl = Moles of NaCl / Kilograms of water

Moles of NaCl = 0.342 mol

Kilograms of water = 1.5 kg

Therefore,

Molality of NaCl = (0.342 mol) / (1.5 kg) = 0.228 mol kg¯1

 

Thus:

Δt = i × Kb × m = (2) × (0.512°C kg mol¯1) × (0.228 mol kg¯1) = 0.23°C

 

The boiling point of pure water is 100°C.

Therefore,

t = 100°C + 0.23°C = 100.23°C

t = 100.23°C


Answer: The NaCl solution will boil at 100.23°C